Evaluate the integral, or show why it does not exist. dr (E - 1)2/3 و [
Answer #1

(+)43 no Given if they se days of most da - (very s (2-1)-(0-10%) 3 = 31(0)3-(-1)s] = --Scan+b)da = *(Cax+b)ht, = 361 +1] (h